OpenStudy (anonymous):

Given the equations 2x+43y=1 and y−913x=9, by what factor would you multiply the first equation so that combining the two equations would eliminate x?

OpenStudy (anonymous):

In order to do this we would have to divide 913 by 2. which equates to 456.5. So in turn we can multiply 456.5(2x+43y)=456.5(1) 913x+19629.5y=456.5 We have to multiply positively because our coefficient is negative subsequently canceling both of them out.

OpenStudy (anonymous):

what would that be in a fraction